OpenStudy (anonymous):

Let cos y = .98. Which is csc y? A. .5 B. 5.05 C. 5 D. 5.5

OpenStudy (souvik):

what is sin y?

OpenStudy (anonymous):

i wasnt given that. @souvik

OpenStudy (souvik):

cosec y=1/sin y so we have to find sin y first...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\sin ^{2}y=1-\cos ^{2}y=1-.98^{2}=\left( 1+.98 \right)\left( 1-.98 \right)=1.98*.02\] =.0396 \[\sin y=\pm \sqrt{.0396},\csc y=\pm \sqrt{\frac{ 1 }{.0396 }}\]

OpenStudy (souvik):

well done.......!